OpenStudy (anonymous):

bending stress..what is the formula for flexural stress in a oval cross sectional beam?

OpenStudy (anonymous):

if the moment of inertia for an ellipse is \[I=\frac{ \pi }{ 4 } a b^{3}\] and bending stress is \[\sigma=\frac{ My }{ I }\] and if it were a 3 point bend test, would my formula be \[\sigma=\frac{ 2FL }{ \pi ab ^{2} }\] F is the applied force and L is the distance between 2 supports. F is applied in the middle.